# The Low-Degree Hardness of Finding Large Independent Sets in Sparse Random Hypergraphs

### v2

@@ 2.1 Notation
Let $\mathbb{N}$ denote the set of nonnegative integers. For $n\in\mathbb{N}$ such that $n\geq 1$, we let $[n]\coloneqq\{1,\ldots,n\}$. Throughout this work, we consider $n\to\infty$ and use the asymptotic notation $O(\cdot),\,\Omega(\cdot),\,o(\cdot)$, etc.
@@ 2.2 Local Algorithms on Hypergraphs
For any integer $s\geq 0$ and $v\in V(H)$, let $N_{s}(H,v)$ denote the the rooted hypergraph with root $v$ and its depth-$s$ neighborhood in $H$, and let $|N_{s}(H,v)|$ denote the number of hyeredges in $N_{s}(H,v)$. An $s$-local algorithm for finding an independent set in a hypergraph $H$ is defined by a measurable $s$-local function $g=g(H,v,\mathbf{X})$. Here, the input is a rooted hypergraph $H$ with root $v$ and depth at most $s$, and a vector $\mathbf{X}\in[0,1]^{|V(H)|}$. The output of the function is 00 (‘out’) or $1$ (‘in’) and is invariant under labeled isomorphisms of $N_{s}(H,v)$. We apply an $s$-local function $g$ to a hypergraph $H$ by assigning each vertex $v$ a label $\mathbf{X}_{v}\sim\textsf{Unif}[0,1]$ independently, and then evaluating $g(\cdot)$ on $N_{s}(H,v)$ and the restriction of the labels $\mathbf{X}$ to $V\left(N_{s}(H,v)\right)$. For any $s$-local function $g$, hypergraph $H$, and $\mathbf{X}\in[0,1]^{|V(H)|}$, we require $\{v\,:\,g(H,v,\mathbf{X})=1\}$ is an independent set with probability $1$. A local algorithm is an $s$-local function for some constant $s$.
@@
As mentioned in §1.4, we will consider two kinds of hypertrees in our proof of Theorem 1.5. First, we describe the $r$-uniform Poisson$(d)$ Galton-Watson Hypertree (denoted $\mathbb{T}_{d}^{GW}$) introduced by Pal and Zhu in [pal2021community]. The process generates a rooted $r$-uniform hypertree $(T,o)$ as follows:
  • start with a root vertex $o$ at level 00.
@@
As sparse random hypergraphs are locally hypertree-like, the performance of a local algorithm on $\mathcal{H}_{r}(n,p)$ is determined, up to the first order, by the expectation of the corresponding local function $g$ evaluated at the root of $\mathbb{T}_{d}^{GW}$. More precisely, we have the following lemma:
@@
Suppose $g$ is an $s$-local function for independent sets. Let $I$ be the random independent set obtained by applying $g$ to $H\sim\mathcal{H}_{r}(n,p)$. Then for $T\sim\mathbb{T}_{d}^{GW}$ with high probability over the randomness in $H$ and the random label $\mathbf{X}\sim\mathsf{Unif}[0,1]^{n}$,
$\frac{|I|}{n}=\mathbb{E}[g(T,o,\mathbf{X})]+o(1).$
@@
The expectation of $|I|$ is $\mathbb{E}[g(T,o,\mathbf{X})]n+o(n)$. Furthermore, the variance of $|I|$ is $O(n)$ as there are at most $O(n)$ pairs of vertices of $u,v$ such that $g(H,u,\mathbf{X})$ and $g(H,v,\mathbf{X})$ are not independent. The lemma follows by Chebyshev’s Inequality. ∎
To assist with our proofs, we define
[…]
@@
The second hypertree we consider is the the rooted $r$-uniform $\Delta$-regular hypertree (denoted $T_{\Delta}^{r}$) with root $o$. Observe that when applying an $s$-local algorithm $h$ to $T_{\Delta}^{r}$ with vertex labels $\mathbf{X}$, the distribution of $h(T_{\Delta}^{r},v,\mathbf{X})$ does not depend on the vertex $v$. Hence we can define the density of the independent set $I$ obtained by applying $h$ to $T_{\Delta}^{r}$ as $density(I)=\mathbb{P}[h(T_{\Delta}^{r},o,\mathbf{X})=1]=\mathbb{E}[h(T_{\Delta}^{r},o,\mathbf{X})]$. Therefore, we can define the following parameter similar to (1):
[…]
@@ 2.3 Interpolating Paths and Stable Algorithms
Let $T\in\mathbb{N}$ and let $m=\binom{n}{r}$ (resp. $m=n^{r}$). We construct a length $T$ interpolation path $A^{(0)},\ldots,A^{(T)}$ as follows: $A^{(0)}\sim\mathcal{H}_{r}(n,p)$ (resp. $\mathcal{H}(r,n,p)$); then for each $1\leq t\leq T$, $A^{(t)}$ is obtained by resampling the coordinate $\sigma(t)\in[m]$ from $Ber\left(d/{n-1\choose r-1}\right)$ (resp. $Ber\left(d/n^{r-1}\right)$). Here $\sigma(t)=t-k_{t}m$ where $k_{t}$ is the unique integer making $1\leq\sigma(t)\leq m$.
Consider the hypercube $Q_{m}$ with vertex set $\{0,1\}^{m}$ and edges of the form $uv$ where $u$ and $v$ differ in exactly one coordinate. Interpolation paths can therefore be viewed as random walks on $Q_{m}$ (for the appropriate value of $m$). We define $(D,\Gamma,c)$-stable algorithms with respect to this random walk (the distribution $\mathcal{P}$ below is either $\mathcal{H}_{r}(n,p)$ or $\mathcal{H}(r,n,p)$).
@@
[no edge of $\mathbb{P}\left[\text{no edge of }A^{(0)},\ldots,A^{(T)}\text{ is $c$-bad for $f$}\right]\geq p^{4\Gamma D/c}.$
The following lemma will be crucial to proving our impossibility results. The proof follows an identical argument to that of [wein2022optimal, Lemmas 2.7 and 2.8] on graphs and so we omit it here.
@@ 3.1.1 Local Algorithms: from $T_{\Delta}^{r}$ to $\mathbb{T}_{d}^{GW}$
For any $\varepsilon>0$ and any sufficiently large $d=d(\varepsilon)>0$, there exists $s\coloneqq s(\varepsilon,d)$ and an $s$-local function $g$ such that for $(T,o)\sim\mathbb{T}_{d}^{GW}$ and i.i.d $\mathsf{Unif}[0,1]$ labels $\mathbf{X}$ over its vertices, $\mathbb{E}[g(T,o,\mathbf{X})]\geq(1-\varepsilon)\left(\frac{1}{r-1}\cdot\frac{\log d}{d}\right)^{\frac{1}{r-1}}$.
We start by describing the following random greedy algorithm analyzed in [nie2021randomized], which generates an independent set in an $r$-uniform $\Delta$-regular hypergraph $H$ with large girth (which implies the case of $T_{\Delta}^{r}$):
@@
We remark that while this algorithm as described is not local, it can be simulated by one that is (this can be inferred by the discussion in [nie2021randomized, §2]). Nie and Verstraëte [nie2021randomized, Theorem 4] showed that the random greedy algorithm outputs an independent set of density $(1-\varepsilon)\left(\frac{1}{r-1}\cdot\frac{\log\Delta}{\Delta}\right)^{1/r-1}$ on $T_{\Delta}^{r}$.
Following an idea of Rahman and Virág in [rahman2017local], we will show that a local algorithm finding an independent set in $T_{\Delta}^{r}$ can be adapted to construct an independent set in $\mathbb{T}_{d}^{GW}$ as well. We remark that the parameter $d$ we choose will not be equal to $\Delta$, however, it will be chosen such that $\log d/d=(1\pm o_{d}(1))\log\Delta/\Delta$. In particular, the independent set output by our local algorithm on $\mathbb{T}_{d}^{GW}$ will have the same density as the one in $T_{\Delta}^{r}$ for $d$ sufficiently large. Let $E(d,\Delta)$ denote the event that the root of $\mathbb{T}_{d}^{GW}$ and all of its neighbors have degree at most $\Delta$. With this definition in hand, we prove the following proposition:
@@
Let $I$ be an independent set generated by $h$ on $T_{\Delta}^{r}$ and some i.i.d $\mathsf{Unif}[0,1]$ vertex labels, and let $T\sim\mathbb{T}_{d}^{GW}$. The local algorithm $g$, that can generate the independent set $J$ of the desired density in $T$, is described explicitly through three steps.
@@ 2. Step 2:
    Regularization. We describe how to construct a $\Delta$-regular hyperforest $T^{\prime}$ from $S$ such that $S\subseteq T^{\prime}$. Define the $(\Delta-1)$-ary hypertree to be the rooted hypertree in which every vertex has exactly $\Delta-1$ offspring edges. Consider $r-1$ such $(\Delta-1)$-ary hypertrees with roots $u_{1},u_{2},\ldots,u_{r-1}$ and create an edge $\{v,u_{1},\ldots,u_{r-1}\}$. Repeat this process $\Delta-\deg_{S}(v)$ many times for each $v\in V(S)$. Randomly label $T^{{}^{\prime}}$ by a new labeling $\mathbf{X}^{\prime}$ independent of $\mathbf{X}$.
@@ 3. Step 3:
    Inclusion. Since $T^{\prime}$ is a disjoint collection of $r$-uniform $\Delta$-regular hypertrees, we can use $h$ with input $\mathbf{X}^{\prime}$ to construct an independent set $I^{\prime}$ of $T^{\prime}$ with the same density $I$. However, due to the removal of edges from $T$, $I^{\prime}$ may no longer be independent in $T$. We construct the desired set $J$ from $I^{\prime}$ by including all vertices $v\in I^{\prime}$ such that no edges containing $v$ were removed during Step 1.
By design, $h$ is a local algorithm on $\mathbb{T}_{d}^{GW}$. For $e\in E(T)$, if $e\subseteq J$, then $e\subseteq I^{\prime}$ as well. Since $I^{\prime}$ is an independent set in $T^{\prime}$, $e$ must have been removed during Step 1. By construction, none of the vertices in $e$ would have been added to $J$. Therefore, $J$ is indeed an independent set in $T$.
Let us now prove the desired bounds on the density of $J$. Clearly, $\mathsf{density}(J)\leq\mathsf{density}(I^{\prime})$ since $J\subseteq I^{\prime}$. Furthermore, for any $v\in I^{\prime}$, if $v$ and all of it’s neighbors in $T$ have degree at most $\Delta$ then none of the edges containing $v$ will be removed, which implies that $v\in J$. Hence the lower bound follows by the previous observation. ∎
@@
If $d=\Delta-\Delta^{u}$ for any $11/2<u<$ 1 then $\mathbb{P}[E(d,\Delta)]\to 1$ as $\Delta\to\infty$.
@@
Let $X$ denote the degree of the root of a $\mathbb{T}_{d}^{GW}$ Tree. Let $\{Z_{i,j}\}_{i\in[X],j\in[r-1]}$ denote the number of offspring edges adjacent to the $j$-th vertex in the $i$-th offspring edge of the root. Recall that $X$ has distribution $\mathsf{Pois}(d)$ and that conditioned on $X$ all $Z_{i,j}$ are i.i.d with distribution $\mathsf{Pois}(d)$. Let $p(d,\Delta)\coloneqq\mathbb{P}[\mathsf{Pois}(d)>\Delta]$. We have
$\mathbb{P}[E(d,\Delta)]$ $=\mathbb{E}\left[\mathbbm{1}_{X\leq\Delta}\cdot\prod_{i=1}^{X}\prod_{j=1}^{r-1}\mathbbm{1}_{Z_{i,j}\leq\Delta-1}\right]$
$=\mathbb{E}\left[\mathbbm{1}_{X\leq\Delta}\cdot\mathbb{E}\left[\prod_{i=1}^{X}\prod_{j=1}^{r-1}\mathbbm{1}_{Z_{i,j}\leq\Delta-1}\Big{|}X\right]\right]$
$=\mathbb{E}\left[\mathbbm{1}_{X\leq\Delta}\cdot\Big{(}1-p(d,\Delta-1)\Big{)}^{(r-1)X}\right]$
$=\mathbb{E}\left[\Big{(}1-p(d,\Delta-1)\Big{)}^{(r-1)X}\right]-\mathbb{E}\left[\mathbbm{1}_{X>\Delta}\cdot\Big{(}1-p(d,\Delta-1)\Big{)}^{(r-1)X}\right].$
Using the moment generating function of a $\mathsf{Pois}(d)$ random variable we have
$\mathbb{E}\left[\Big{(}1-p(d,\Delta-1)\Big{)}^{(r-1)X}\right]$ $=\mathbb{E}\left[\exp\Big{(}(r-1)X\log\big{(}1-p(d,\Delta-1)\big{)}\Big{)}\right]$
$=\exp\left(d\left(e^{(r-1)\log\big{(}1-p(d,\Delta-1)\big{)}}-1\right)\right)$
$=\exp\left(d\left(\big{(}1-p(d,\Delta-1)\big{)}^{(r-1)}-1\right)\right)$
$\geq\exp\left(d\Big{(}1-(r-1)p(d,\Delta-1)-1\Big{)}\right)$
$=\exp\left(-d(r-1)p(d,\Delta-1)\right)$
Additionally,
$\mathbb{E}\left[\mathbbm{1}_{X>\Delta}\cdot\Big{(}1-p(d,\Delta-1)\Big{)}^{(r-1)X}\right]$ $=\sum_{x>\Delta}\Big{(}1-p(d,\Delta-1)\Big{)}^{(r-1)x}\cdot\mathbb{P}[X=x]$
$\leq\Big{(}1-p(d,\Delta-1)\Big{)}^{\Delta(r-1)}\sum_{x>\Delta}\mathbb{P}[X=x]$
$=\Big{(}1-p(d,\Delta-1)\Big{)}^{\Delta(r-1)}\cdot p(d,\Delta)$
$\leq p(d,\Delta)\leq p(d,\Delta-1)$
Putting the above together, we have
$\mathbb{P}[E(d,\Delta)]\geq\exp\left(-d(r-1)p(d,\Delta-1)\right)-p(d,\Delta-1).$
From here, we may complete the proof by an exponential moment argument to bound $p(d,\Delta-1)$. The proof follows identically to that in [rahman2017local, Lemma 4.6] and so we omit the details. ∎

### v3

@@ 2.1 Notation
Let $\mathbb{N}$ denote the set of nonnegative integers. For $n\in\mathbb{N}$ such that $n\geq 1$, we let $n]\coloneqq\{1,\ldots,n\}$. Throughout this work, we use the asymptotic notation $O_{x}(\cdot)$, $\Omega_{x}(\cdot)$, $o_{x}(\cdot)$, etc. as $x\to\infty$. We ignore the subscript when $x$ is clear from context. Note that our results are in the double limit $n\to\infty$ followed by $d\to\infty$. In particular, $o_{n}(1)$ is a quantity depending arbitrarily on $d$ tending to 0 as $n\to\infty$, while $o_{d}(1)$ is a quantity independent of $n$ which tends to 0 as $d\to\infty$.
@@ 2.2 Local Algorithms on Hypergraphs
For any integer $s\geq 0$ and $v\in V(H)$, let $N_{s}(H,v)$ denote the the rooted hypergraph with root $v$ and its depth-$s$ neighborhood in $H$, and let $|N_{s}(H,v)|$ denote the number of hyperedges in $N_{s}(H,v)$. An $s$-local algorithm for finding an independent set in a hypergraph $H$ is defined by a measurable $s$-local function $g=g(H,v,\mathbf{X})$. Here, the input is a rooted hypergraph $H$ with root $v$ and depth at most $s$, and a vector $\mathbf{X}\in[0,1]^{|V(H)|}$. The output of the function is 0 (‘out’) or $1$ (‘in’) and is invariant under labeled isomorphisms of $N_{s}(H,v)$. We apply an $s$-local function $g$ to a hypergraph $H$ by assigning each vertex $v$ a label $\mathbf{X}_{v}\sim\textsf{Unif}[0,1]$ independently, and then evaluating $g(\cdot)$ on $N_{s}(H,v)$ and the restriction of the labels $\mathbf{X}$ to $V\left(N_{s}(H,v)\right)$. We say an $s$-local function $g$ is valid if for any hypergraph $H$, and any label $\mathbf{X}\in[0,1]^{|V(H)|}$, the set $\{v\,:\,g(H,v,\mathbf{X})=1\}$ is an independent set in $H$. A local algorithm is a valid $s$-local function for some constant $s$.
@@
As mentioned in §1.4, we will consider two kinds of hypertrees in our proof of Theorem 1.5. First, we describe the $r$-uniform Poisson(d)(d) Galton-Watson Hypertree (denoted $\mathbb{T}_{d}^{GW}$) introduced by Pal and Zhu in [PZ21]. The process generates a rooted $r$-uniform hypertree $(T,o)$ as follows:
  • start with a root vertex $o$ at level 0.
@@
It is a well-known fact that sparse random graphs are locally tree-like. The same holds for hypegraphs [^3] To the best of our knowledge, no proof of this fact exists in the literature, however, the proof is nearly identical to the graph case. For completeness, we include such a proof in §A. and so the performance of a local algorithm on $\mathcal{H}_{r}(n,p)$ is determined, up to the first order, by the expectation of the corresponding local function $g$ evaluated at the root of $\mathbb{T}_{d}^{GW}$. More precisely, we have the following lemma:
@@
Suppose $g$ is a valid $s$-local function for independent sets. Let $I$ be the random independent set obtained by applying $g$ to $H\sim\mathcal{H}_{r}(n,p)$. Then for $T\sim\mathbb{T}_{d}^{GW}$ with high probability over the randomness in $H$ and the random label $\mathbf{X}\sim\mathsf{Unif}[0,1]^{n}$,
$\frac{|I|}{n}=\mathbb{E}[g(T,o,\mathbf{X})]+o_{n}(1).$
To assist with the proof of the above result, we will employ the following proposition.
Proposition 2.2.
Fix some integer $s$ and $r\geq 2$. Let $H\sim\mathcal{H}_{r}(n,p)$ for $p=d/{n-1\choose r-1}$ where $d\geq 1$ and let $\mathbf{X}\sim\mathsf{Unif}[0,1]^{n}$. Let $g$ be an $s$-local function taking values in $\{0,1\}$. We have
$\mathbb{E}\left[\bigg|\sum_{v\in[n]}g(H,v,\mathbf{X})-\mathbb{E}\sum_{v\in[n]}g(H,v,\mathbf{X})\bigg|^{2}\right]=O(n)$
@@
Fix some order on the vertices of $H$. Let $E_{x}$ be set $\{e\in K_{n}^{r}\,:\,x=\min e\}$. Note that the random variables $(E_{x})_{x\in[n]}$ are independent. Furthermore, there is a function $F$ such that
$\sum_{v\in[n]}g(H,v,\mathbf{X})=F(E_{1},\ldots,E_{n}).$
Let $Y$ denote this sum and let $Y_{x}=F(E_{1},\ldots,E_{x-1},\emptyset,E_{x+1},\ldots,E_{n})$.
Equivalently, if $H_{x}$ indicates the subhypergraph of $H$ where all hyperedges containing $x$ as the minimal vertex have been deleted, we have
$Y_{x}=\sum_{v\in[n]}g(H_{x},v,\mathbf{X}).$
Since $(E_{x})_{x\in[n]}$ are independent and $0\leq g(H,v,\mathbf{X})\leq 1$ for all $v$, we can apply the moment inequality [BLM13, Theorem 15.5]: there exists some constant $c_{1}>0$ such that
$\mathbb{E}\left[\Bigg|\sum_{v}g(H,v,\mathbf{X})-\mathbb{E}\sum_{v}g(H,v,\mathbf{X})\Bigg|^{2}\right]\leq c_{1}\mathbb{E}\left[\sum_{x}(Y-Y_{x})^{2}\right].$
Now fix vertices $v$ and $x$ of $H$. Since $g$ is $s$-local, we must have $g(H,v,\mathbf{X})-g(H_{x},v,\mathbf{X})=0$ unless $v$ is in the $s$-neighborhood of $x$, denoted $N_{s}(H,x)$. This implies
$|Y-Y_{x}|=\left|\sum_{v\in N_{s}(H,x)}g(H,v,\mathbf{X})-g(H_{x},v,\mathbf{X})\right|\leq\sum_{v\in N_{s}(H,x)}g(H,v,\mathbf{X})+g(H_{x},v,\mathbf{X})\leq 2\left|N_{s}(G,x)\right|.$
It follows that
$\mathbb{E}\left[\sum_{x}(Y-Y_{x})^{2}\right]\leq 4\mathbb{E}\left[\sum_{x}\left|N_{s}(H,x)\right|^{2}\right]=4n\mathbb{E}\left[\left|N_{s}(H,x)\right|^{2}\right]$
Note that $|N_{s}(H,x)|\mid|N_{s-1}(H,x)|\sim(r-1)|N_{s-1}(H,x)|\mathsf{Pois}(d)$. By repeated applications of the tower property of expectation, one can show that there exists some constant $c_{2}\coloneqq c_{2}(r,s)>0$ such that
$\mathbb{E}\left[\left|N_{s}(H,x)\right|^{2}\right]\leq c_{2}d^{2s}.$
Putting everything together, we have
$\mathbb{E}\left[\Bigg|\sum_{v\in R}g_{r}(H_{v},\mathbf{x})-\mathbb{E}\sum_{v\in R}g_{r}(H_{v},\mathbf{x})\Bigg|^{2}\right]\leq 4c_{1}c_{2}d^{2s}n,$
as desired. ∎
We note that Wein employed a version of the above result for graphs [Wei22, Corollary 3.5], which he obtained as a corollary to a special case of a more general statement on graph functionals [BCN23, Proposition 12.3]. With this in hand, we are ready to prove Lemma 2.1.
Proof of Lemma 2.1.
Utilizing Proposition A.1, the local weak convergence of $\mathcal{H}_{r}(n,p)$ to $\mathbb{T}_{d}^{GW}$, it is easy to deduce that the expectation of $|I|$ is $\mathbb{E}[g(T,o,\mathbf{X})]n+o(n)$. Furthermore, Proposition 2.2 implies that the variance is $O(n)$. The claim now follows by Chebyshev’s inequality. ∎
Observe that when applying an $s$-local algorithm $g$ to $T\sim\mathbb{T}_{d}^{GW}$ with vertex labels $\mathbf{X}$, the distribution of $g(T,v,\mathbf{X})$ does not depend on the vertex $v$, since the distribution of $N_{s}(T,v)$ does not depend on the choice of $v$ in $T$. Hence we can define the density of the independent set $I$ obtained by applying $g$ to $T$ as $\mathsf{density}(I)=\mathbb{P}[h(T,o,\mathbf{X})=1]=\mathbb{E}[h(T,o,\mathbf{X})]$. To assist with our proofs, we define
$\alpha_{r}(d)\coloneqq\left(\frac{d}{\log d}\right)^{\frac{1}{r-1}}\,\sup\left\{\mathbb{E}[g(T,o,\mathbf{X})]\,:\,s\geq 0,\,g\text{ is a valid }s\text{-local function}\right\}.$ (1)
@@
The second hypertree we consider is the the rooted $r$-uniform $\Delta$-regular hypertree (denoted $T_{\Delta}^{r}$) with root $o$. Again, observe that when applying an $s$-local algorithm $h$ to $T_{\Delta}^{r}$ with vertex labels $\mathbf{X}$, the distribution of $h(T_{\Delta}^{r},v,\mathbf{X})$ does not depend on the vertex $v$. Hence we can define the density of the independent set $I$ obtained by applying $h$ to $T_{\Delta}^{r}$ as $\mathsf{density}(I)=\mathbb{P}[h(T_{\Delta}^{r},o,\mathbf{X})=1]=\mathbb{E}[h(T_{\Delta}^{r},o,\mathbf{X})]$. Therefore, we can define the following parameter similar to (1):
$\alpha(r,\Delta)\coloneqq\left(\frac{\Delta}{\log\Delta}\right)^{\frac{1}{r-1}}\,\sup\left\{\mathbb{E}[h(T_{\Delta}^{r},o,\mathbf{X})]\,:\,s\geq 0,\,h\text{ is a valid }s\text{-local function}\right\}.$ (2)
@@ 2.3 Interpolating Paths and Stable Algorithms
Let $T\in\mathbb{N}$ and let $m=\binom{n}{r}$ (resp. $m=n^{r}$). We construct a length $T$ interpolation path $A^{(0)},\ldots,A^{(T)}$ as follows: $A^{(0)}\sim\mathcal{H}_{r}(n,p)$ (resp. $\mathcal{H}(r,n,p)$); then for each $1\leq t\leq T$, $A^{(t)}$ is obtained from $A^{(t-1)}$ by resampling the coordinate $\sigma(t)\in[m]$ from $Ber\left(d/{n-1\choose r-1}\right)$ (resp. $Ber\left(d/n^{r-1}\right)$). Here $\sigma(t)=t-k_{t}m$ where $k_{t}$ is the unique integer satisfying $1\leq\sigma(t)\leq m$.
Consider the hypercube $Q_{m}$ with vertex set $\{0,1\}^{m}$ and edges of the form $uv$ where $u$ and $v$ differ in exactly one coordinate. Interpolation paths can therefore be viewed as random walks on $Q_{m}$ (for the appropriate value of $m$). We define $(D,\Gamma,c)$-stable algorithms with respect to this random walk (the distribution $\mathcal{P}$ below is either $\mathcal{H}_{r}(n,p)$ or $\mathcal{H}(r,n,p)$ with the edge probability $p$ set to $d/{n-1\choose r-1}$ or $d/n^{r-1}$, respectively).
@@
$\mathbb{P}\left[\text{no edge of }A^{(0)},\ldots,A^{(T)}\text{ is $c$-bad for $f$}\right]\geq p^{4\Gamma D/c}.$
The following lemma will be crucial to proving our impossibility results. The proof follows an identical argument to that of [Wei22, Lemmas 2.7 and 2.8] on graphs and so we omit it here.
@@ 3.1.1 Local Algorithms: from $T_{\Delta}^{r}$ to $\mathbb{T}_{d}^{GW}$
For any $\varepsilon>0$ and any sufficiently large $d=d(\varepsilon)>0$, there exists $s\coloneqq s(\varepsilon,d)$ and a valid $s$-local function $g$ such that for $(T,o)\sim\mathbb{T}_{d}^{GW}$ and i.i.d $\mathsf{Unif}[0,1]$ labels $\mathbf{X}$ over its vertices, $\mathbb{E}[g(T,o,\mathbf{X})]\geq(1-\varepsilon)\left(\frac{1}{r-1}\cdot\frac{\log d}{d}\right)^{\frac{1}{r-1}}$.
We start by describing the following random greedy algorithm analyzed in [NV21], which generates an independent set in an $r$-uniform $\Delta$-regular hypergraph $H$ with large girth (which implies the case of $T_{\Delta}^{r}$):
@@
We remark that while this algorithm as described is not local, it can be simulated by one that is (this can be inferred by the discussion in [NV21, §2]). Nie and Verstraëte [NV21, Theorem 4] showed that the random greedy algorithm outputs an independent set of density $(1-\varepsilon)\left(\frac{1}{r-1}\cdot\frac{\log\Delta}{\Delta}\right)^{1/r-1}$ on $T_{\Delta}^{r}$.
Following an idea of Rahman and Virág in [RV17], we will show that a local algorithm finding an independent set in $T_{\Delta}^{r}$ can be adapted to construct an independent set in $\mathbb{T}_{d}^{GW}$ as well. We remark that the parameter $d$ we choose will not be equal to $\Delta$, however, it will be chosen such that $\log d/d=(1\pm o_{d}(1))\log\Delta/\Delta$. In particular, the independent set output by our local algorithm on $\mathbb{T}_{d}^{GW}$ will have the same density as the one in $T_{\Delta}^{r}$ for $d$ sufficiently large. Let $E(d,\Delta)$ denote the event that the root of $\mathbb{T}_{d}^{GW}$ and all of its neighbors have degree at most $\Delta$. With this definition in hand, we prove the following proposition:
@@
Let $I$ be an independent set generated by $h$ on $T_{\Delta}^{r}$ and some i.i.d $\mathsf{Unif}[0,1]$ vertex labels, and let $T\sim\mathbb{T}_{d}^{GW}$. The local algorithm $g$, that can generate the independent set $J$ of the desired density in $T$, is described explicitly through the following three steps.
@@ 2. Step 2:
    Regularization. We describe how to construct a $\Delta$-regular hyperforest $T^{\prime}$ from $S$ such that $S\subseteq T^{\prime}$. Define the $(\Delta-1)$-ary hypertree to be the rooted hypertree in which every vertex has exactly $\Delta-1$ offspring edges. Consider r1r-1 such $(\Delta-1)$-ary hypertrees with roots $u_{1},u_{2},\ldots,u_{r-1}$ and create an edge $\{v,u_{1},\ldots,u_{r-1}\}$. Repeat this process $\Delta-\deg_{S}(v)$ many times for each $v\in V(S)$. Randomly label $T^{{}^{\prime}}$ by a new labeling $\mathbf{X}^{\prime}$ independent of $\mathbf{X}$.
@@ 3. Step 3:
    Inclusion. Since $T^{\prime}$ is a disjoint collection of $r$-uniform $\Delta$-regular hypertrees, we can use $h$ with input $\mathbf{X}^{\prime}$ to construct an independent set $I^{\prime}$ of $T^{\prime}$ with the same density as $I$. However, due to the removal of edges from $T$, $I^{\prime}$ may no longer be independent in $T$. We construct the desired set $J$ from $I^{\prime}$ by including all vertices $v\in I^{\prime}$ such that no edges containing $v$ were removed during Step 1.
By design, $g$ is a local algorithm on $\mathbb{T}_{d}^{GW}$. We now show that $J$ is an independent set. Suppose for contradiction that $e\subseteq J$ for some $e\in E(T)$. Then $e\subseteq I^{\prime}$ as well. Since $I^{\prime}$ is an independent set in $T^{\prime}$, $e$ must have been removed during Step 1. By the construction in the last step, none of the vertices in $e$ would have been added to $J$. Therefore, $J$ is indeed an independent set in $T$.
Let us now prove the desired bounds on the density of $J$. Clearly, $\mathsf{density}(J)\leq\mathsf{density}(I^{\prime})$ since $J\subseteq I^{\prime}$. Furthermore, for any $v\in I^{\prime}$, if $v$ and all of its neighbors in $T$ have degree at most $\Delta$ then none of the edges containing $v$ will be removed, which implies that $v\in J$. Hence the lower bound follows by the previous observation. ∎
@@
If $d=\Delta-\Delta^{u}$ for any $1/2<u<1$ then $\mathbb{P}[E(d,\Delta)]\to 1$ as $\Delta\to\infty$.
@@
Let $p(d,\Delta)\coloneqq\mathbb{P}[\mathsf{Pois}(d)\geq\Delta]$. A direct Chernoff bound using the moment generating function of $\mathsf{Pois}(d)$ gives
$p(d,\Delta)\leq p_{err}\coloneqq\exp\left\{-\Big(1+O\left(\Delta^{u-1}\right)\Big)\frac{\Delta^{2u-1}}{2}\right\}.$
By a simply union bound, with probability at least $1-r\Delta\cdot p_{err}$, the event $E(d,\Delta)$ holds. ∎